OpenStudy (alyssa_xo):

Flip a coin until heads shows, assume that the probability of heads on one flip is \(\large \frac{3}{5}\). We define a RV (random variable) X = number of flips. Find the probability that X will be no more than 10.

OpenStudy (alyssa_xo):

So I have the geometric series \(\huge \frac{3}{5}*\frac{2}{5}^n\). I assume the numerator of the probability goes from 0 (or 1?) to 10

OpenStudy (alyssa_xo):

\[\huge \frac{3}{5}*\frac{2}{5}^n\]

OpenStudy (alyssa_xo):

is latex broked

OpenStudy (anonymous):

i think it is almost 1 \[\frac{3}{5}\sum_{k=0}^{10}(\frac{2}{5})^k\]

OpenStudy (anonymous):

first flip \(.6\) second flip \(.4\times .6\) third flip \(.4^2\times .6\) fourth flip \(.4^3\times .6\) etc ok make that \[\frac{3}{5}\sum_{k=0}^{9}(\frac{2}{5})^k\] when you add